Steam is expanded adibatically in a turbine. If 1000kg/hr of steam at a pressure of 5 bar and 148 degree celcius of superheat enters in the turbine and discharge as saturated steam at 1 bar , calculat the work output of the turbine.


1
Expert's answer
2020-12-24T03:53:22-0500

1000kg/hr of steam at a pressure of 5 bar 


148 degree celcius


1000kg = 1 hr


(P1V1)/T1 = (P2V2)/T2

((1000)(5))/441 = 1(V2)/ 441


V2 / 441 = 5000/441


V2 = (5000/441)(441)


V2 = 5000kg